The equation of a curve is y=tan(x+1). Find, to two decimal places, the gradient of the normal to the curve at the point where x=0.
Answer by Alan3354(69443)

The equation of a curve is y=tan(x+1). Find, to two decimal places, the gradient of the normal to the curve at the point where x=0.
--------
y = tan(x+1)
y' = sec^2(x+1) - 1st derivative
y'(0) = sec^2(1) = 3.4255
--> -1/3.4255 =~ -0.29